OpenStudy (anonymous):

Prove the following: If p is prime and n is any positive integer, then sum of divisors of (p^n) is (p^(n+1)-1)/(p-1). Prove the following: If p is prime and n is any positive integer, then sum of divisors of (p^n) is (p^(n+1)-1)/(p-1). @Mathematics

OpenStudy (kinggeorge):

Are we including 1 and p^n?

OpenStudy (kinggeorge):

You have to do a proof by induction for this one. First, we need to state that the sum of all the divisors of \[p^n\] are \[1+ p+ p^2+ p^3+ ...+ p^n\] since p is prime. Then we start the proof by induction. First, let \[n = 0\] Then \[p^0 = 1 = {p^{0+1} - 1 \over p - 1}\] Then we assume this is true for some \[k \in \mathbb{N}\] so \[1+ p + p^2+...+p^k = {p^{k+1} - 1 \over p - 1}\] Then\[1+p+...p^k+p^{k+1} = {p^{k+1} - 1 \over p - 1} + p^{k+1} = {(p^{k+1} - 1) + p^{k+1}(p - 1) \over p - 1}\] So \[1+p+...+p^{k+1} = {p^{(k+1)+1} + p^{k+1}-p^{k+1} -1 \over p -1} = {p^{(k+1)+1} - 1 \over p-1}\] Which is what we wanted so therefore, the sum of the divisors of \[p^n\] where p is prime, is \[{p^{n+1} - 1 \over p -1} \forall n \in \mathbb{N}\]

OpenStudy (anonymous):

trying to use this one to prove the other one I asked before. I can't justify how the sum of the divisors 2^(k-1) is (2^(k+1)-1)

OpenStudy (kinggeorge):

Where do you need the sum of the divisors of 2^(k-1)?

OpenStudy (anonymous):

http://primes.utm.edu/notes/proofs/EvenPerfect.html in this proof

OpenStudy (kinggeorge):

Wouldn't you be trying to show the sum of the divisors of 2^(k-1) is ((2^k) -1)?

OpenStudy (anonymous):

I can easily say that the sum of the divisors of a prime (2^k)-1 is just 2^k since it's just 1+ (2^k)-1 but the sum of the divisors of 2^(k-1) is a lot harder to show

OpenStudy (anonymous):

yes you are right sorry

OpenStudy (kinggeorge):

Well, shouldn't that be almost trivial now since 2 is prime? The factors of 2^(k-1) should sum to \[{2^k -1 \over 2-1} = 2^k -1\] by what we just proved.

OpenStudy (anonymous):

the problem with this way is that it's assuming that I know what the form of the sum should be in and almost proving it backwards. It's like I'm assuming something I shouldn't know. Is there a simpler way to show that 2^(k-1) is ((2^k) -1)?

OpenStudy (anonymous):

maybe a way to show that it equals this in this specific case

OpenStudy (kinggeorge):

To the best of my knowledge, that's the best way to do it.

OpenStudy (kinggeorge):

We could do another proof by induction using 2 instead of p, but that's just a degenerate case of what we proved earlier.

OpenStudy (anonymous):

the thing is that I'm supposed to build the proofs off of theorems and propositions I already know. So I can't just go and prove it by knowing the conclusion. The reason why I asked this question was so I could understand where it was coming from better so I could maybe derive it myself without first knowing the form of the conclusion

OpenStudy (kinggeorge):

You could try making some sort of combinatorial argument, but I'm not sure how well it work. I can't help you much more.

OpenStudy (anonymous):

can you prove this by induction? If k is a positive integer and (2^k)-1 is prime, then (2^(k-1))*((2^k)-1) is perfect.

OpenStudy (anonymous):

I was think that it might be possible to do it that way since the conclusion is would be a sum of numbers which is a form that induction would be good for.

OpenStudy (kinggeorge):

I don't think that can be proven by induction because 2^(k+1) - 1 is not necessarily prime.

OpenStudy (kinggeorge):

Good luck!

OpenStudy (anonymous):

thanks for your help, I will do my best!